OpenStudy (anonymous):

HELP ME PLEASE !!! What is the equation of the line that passes through the points (4/5 , 1/5) and1/2 , 3/2 ? A. y = –x – 1 B. y = –x + 1 C. Y=-13/3X+11/3 D. Y=7/3X+1/3

OpenStudy (anonymous):

PLEASE HELP ME

OpenStudy (anonymous):

Do you want the entire work or just the answer?

OpenStudy (anonymous):

Try it m=y2-y1/x2-x1

OpenStudy (anonymous):

Per exemple x1=4/5 and x2=1/2;

OpenStudy (anonymous):

whats the answer in on a time limit :(

OpenStudy (anonymous):

letter C

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

Plz gimme a medal if the work is worth.